• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Camsonne

    Log entry time 08:34:07 on August 18, 2006

    Entry number 180619

    keyword=Owl shift summary

    Run summary LEDEX

    Owl shift summary

    08/18/2006

    LEDEX gammaD Kin point 25



    Shift crew :
    A. Camsonne
    Michael Olson
    Lee Byungwuek




    0:00 No beam... MCC working on optics

    2:53 Beam in Hall A

    3:10 Optics run 7 foil done

    3:15  moved to  4 cm dummy target

    3:29 moved to  15 cm dummy target

    3:43 moved to  Tantalu target

    3:57 moved to  LD2 for production with 5% radiator

    Current runs seem to have very few FPP hits, called Jackie and Mark they are looking into it.

    All seem fine to me as far as FPP is concerned (HV,LV, threshold, Gas Flow )
    continuing data taking this way




    Run list

    Run Target Rad N T3(M) Total (M) Replay Right Arm
    2854 C optics
    out


    OK 21197
    2855 Dummy 4 cm
    out


    OK 21198
    2856 Dummy 15
    out


    OK 21199
    2857 Tantalum
    out


    OK 21200







    2858 LD2 5% 1.9 39.7
    FPP low efficiency
    21201
    2859 LD2 5% 1.9 40.6
    FPP low efficiency 21202
    2860 LD2 5% 1.9 42.5
    FPP low efficiency 21203
    2861 LD2 5% 1.9 44.4
    FPP low efficiency 21204
    2862 LD2 5% 1.9 46.3
    FPP low efficiency 21205
    2863 LD2 5% 1.9 48.2
    FPP low efficiency 21206
    2864 LD2 5% 1.9 50.1
    FPP low efficiency 21207
    2865 LD2 5% 1.9 52.0
    FPP low efficiency 21208
    2866 LD2 5% 1.9 53.9
    FPP low efficiency 21209












    Pending issues :
    FPP efficiency seems to be very low might be due to gas ( maybe )